# DenseDPO: Fine-Grained Temporal Preference Optimization for Video Diffusion Models

### v1

@@ Appendix D Limitations and Future Works
Although this work focuses on improving visual quality and consistency while holding motion dynamics constant, the same methodology can be extended to other dimensions. For example, slightly perturbing the text prompt during guided synthesis could support better alignment with textual descriptions. It may even lead to preference labels “for-free", e.g., the video generated with the original prompt is often better than the one generated with the perturbed prompt. Similar approaches have also been studied in [10], where they intentionally distort a video via inversion.
@@ Appendix E Pytorch-style Pseudo Code for StructuralDPO and DenseDPO
␣␣␣␣#model:Flowmodelthattakestextembeddingscandtimestept
␣␣␣␣#␣␣␣␣␣asinputsandpredictsvelocityc
␣␣␣␣#ref_model:Pre-trainedmodelthatisfrozen
␣␣␣␣#x_0:Thefirstvideointhepair,shape[B,T,C,H,W]
␣␣␣␣#x_1:Thesecondvideointhepair,shape[B,T,C,H,W]
␣␣␣␣#c:Textembeddingoftheprompt
␣␣␣␣#l:Preferencelabel,shape[B],eachitemiseither+1or-1
␣␣␣␣#␣␣␣␣␣+1meansx_0isbetterthanx_1,-1meanstheotherway
␣␣␣␣#beta:DPOregularizationhyper-parameter
␣␣␣␣#returns:Flow-DPOlossvalue
@@
␣␣␣␣#model:Flowmodelthattakestextembeddingscandtimestept
␣␣␣␣#␣␣␣␣␣asinputsandpredictsvelocityc
␣␣␣␣#ref_model:Pre-trainedmodelthatisfrozen
␣␣␣␣#x_0:Thefirstvideointhepair,shape[B,T,C,H,W]
␣␣␣␣#x_1:Thesecondvideointhepair,shape[B,T,C,H,W]
␣␣␣␣#c:Textembeddingoftheprompt
␣␣␣␣#l:Densepreferencelabel,shape[B,T],canbe+1,0,or-1
␣␣␣␣#␣␣␣␣␣+1meansx_0isbetterthanx_1,-1meanstheotherway
␣␣␣␣#␣␣␣␣␣0meansatie
␣␣␣␣#beta:DPOregularizationhyper-parameter
␣␣␣␣#returns:Flow-DPOlossvalue

### v2

@@ Appendix D Limitations and Future Works
Finally, our segment-level preference optimization method can be useful beyond the DPO framework. Our experiments show that even SOTA video reward models are bad at assessing all aspects of a generated video, yet effective online RL training relies on good reward feedback [45, 82]. An interesting direction is to leverage VLM to provide higher-quality dense reward feedback, and adapt RL methods to optimize such dense reward signals.
@@ Appendix E Pytorch-style Pseudo Code for StructuralDPO and DenseDPO
# model: Flow model that takes text embeddings c and timestep t
# as inputs and predicts velocity c
# ref_model: Pre-trained model that is frozen
# x_0: The first video in the pair, shape [B, T, C, H, W]
# x_1: The second video in the pair, shape [B, T, C, H, W]
# c: Text embedding of the prompt
# l: Preference label, shape [B], each item is either +1 or -1
# +1 means x_0 is better than x_1, -1 means the other way
# beta: DPO regularization hyper-parameter
# returns: Flow-DPO loss value
@@
# model: Flow model that takes text embeddings c and timestep t
# as inputs and predicts velocity c
# ref_model: Pre-trained model that is frozen
# x_0: The first video in the pair, shape [B, T, C, H, W]
# x_1: The second video in the pair, shape [B, T, C, H, W]
# c: Text embedding of the prompt
# l: Dense preference label, shape [B, T], can be +1, 0, or -1
# +1 means x_0 is better than x_1, -1 means the other way
# 0 means a tie
# beta: DPO regularization hyper-parameter
# returns: Flow-DPO loss value